How many integers between 999 to 9999 either begin or end with 3 ?

Grade:12th pass

1 Answers

Arun
6 years ago
Let A : the set of integers between 999 and 9999 begin with 3
Let B : the set of integers between 999 and 9999 ends with 3
A = 1 * 103 (begin with 3= 9:1, 9:10, 9:10, 9:10)
B = 9 * 10 * 10 * 1 (end with 3= 9:9, 9:10, 9:10, 9:1)
AB = 1 *102 *1 (begin and end with 3 9:1, 9:10, 9:10, 9:1)
AU B = A + B - AB
= 103 + 900 – 100
= 1000 + 900 – 100
= 1800
8